OpenStudy (anonymous):

The half-life of a certain radioactive material is 37 days. An initial amount of the material has a mass of 477 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 6 days. Round your answer to the nearest thousandth.

OpenStudy (campbell_st):

ok... so the model will be \[P = P_{0} e^{-kt} \] so P is population, Po is initial population, k is the growth/decay constant and t = time you need to find the value of k... using the half life information. so t = 37, Po = 477 and P = 238.5 (half life quantity \[238.5 = 477e^{-37k}\] so \[0.5 = e^{-37t}\] can you solve for t..?

OpenStudy (anonymous):

im going to be honest i have no idea what to do..

OpenStudy (campbell_st):

oops slight typo above ... we need to solve for k... the decay constant the model should read \[0.5 = e^{-37k}\] ok... to solve for k you need to take the base e log of both sides this will remove e so ln(0.5) = 37k so k = -ln(0.5)/37 so k = 0.018734 so you exponential model is \[P = 447e^{-0.018734t}\] all you need to do is substitute t = 6 and evaluate hope this helps

OpenStudy (campbell_st):

to make it easy the model is \[P = 447 \times e^{-0.018734 \times t}\]

OpenStudy (anonymous):

0 kg 426.288 kg 0.736 kg 0.184 kg my answer didnt match any of these i got 399.476?

OpenStudy (campbell_st):

the way to get the correct answer is to use brackets... \[P = 477\times e^{(-0.018734 \times 6)}\]